OpenStudy (ttp):

Hillary has the following scores on her math tests: 65, 86, 95, 78. How much would her mean change if the lowest test score is dropped? (round to the nearest whole number) 5 points 4 points 3 points 2 points

OpenStudy (johnweldon1993):

Well the mean as it is....is 65 + 86 + 95 + 78....and the result of that divided by 4....so 81... if the lowest score (65) is dropped...we will have 86 + 95 + 78 = 259 but now since there are only 3 test...we divide this number by 3... so 259 / 3 = ?

OpenStudy (ttp):

3

OpenStudy (johnweldon1993):

well 259 / 3 = 86.33 so the change from 81 to 86.3 is approximately...?

OpenStudy (ttp):

5?

OpenStudy (johnweldon1993):

Yeah 5...did that make sense? you don't seem to sure

OpenStudy (ttp):

Yes, that made sense. I just double guess myself lol.

OpenStudy (johnweldon1993):

lol just go with your first instinct! :) but good job!

OpenStudy (ttp):

Thanks for helping me! :)

OpenStudy (johnweldon1993):

Anytime! :)